Circular No. 7584

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 WT_{168}

After the publication (MPS 23043) of the initial observations of the apparently asteroidal object 2000 WT_{168} (discovery data below), linkage to further observations (including prediscovery data) showed the orbit to be cometary, although observations did not show cometary activity in December (cf. MPEC 2000-Y21). CCD exposures taken with the 1.5-m reflector at Catalina on 2001 Feb. 13.3 UT by C. W. Hergenrother, however, do show the object to be cometary (highly condensed 9′′.7 coma with red mag 16.3 and 8′′.0 tail in p.a. 110^{o}). Confirmation of cometary activity has been obtained in CCD observations by J. Tichá and M. Tichý at Kleť on Feb. 16.9 (0.57-m reflector; 9′′ tail in p.a. 155^{o} and faint asymmetric coma) and by M. Hicks and B. Buratti at Palomar on Feb. 17.2 (1.5-m reflector; faint teardrop-shaped tail ∼ 15′′ long in p.a. 60^{o}).

2000 UT α_{2000} δ_{2000} m_{2} Observer
Nov. 25.43730 8^{h}35^{m}58.^{s}21 +13^{o}46′16′′.8 17.2 LONEOS
27.37208 8 37 22.96 +13 15 55.8 17.8 LINEAR

The following orbital elements appear on MPEC 2001-D05, which includes additional unpublished astrometry and an ephemeris:

Epoch = 2001 Apr. 1.0 TT

T = 2001 Mar. 23.3229 TT  ω = 245.^{o}4986
e = 0.546608  Ω = 272.5524 }2000.0
q = 1.761729 AU  i = 18.5201

a = 3.885661 AU n^{o} = 0.1286787 P = 7.659 years

COMET C/1999 T1 (McNAUGHT-HARTLEY) Ephemeris from orbital elements on MPC 42106:

2001 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Feb. 10 16^{h}47.^{m}57 +23^{o}34.^{′}5 1.297 1.483 79.^{o}8 40.^{o}9  8.4
20 17 14.12 +33 13.4 1.343 1.578 83.7 38.5  8.7
Mar. 2 17 39.23 +41 57.7 1.424 1.679 86.1 36.1  9.1
12 18 02.25 +49 31.4 1.532 1.784 87.1 33.8  9.4
22 18 22.44 +55 54.1 1.659 1.892 87.1 31.7  9.8
Apr. 1 18 38.91 +61 13.2 1.799 2.002 86.4 29.9 10.2
11 18 50.62 +65 37.6 1.944 2.113 85.3 28.2 10.5
21 18 56.35 +69 15.7 2.092 2.225 84.1 26.7 10.9
2001 February 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green